I keep on getting the wrong answer. I am recording a 105,937.50 for the 2nd year. the answer says $129,479.

Would I be able to get a thorough step by step answer in order for me to see what I am doing wrong with the Double Declining Balance Method. Thank you.

  1. Beauty Company purchased a machine valued at $565,000 on September 1. The equipment has an estimated useful life of eight years or 5.5 million units. The equipment is estimated to have a salvage value of $48,300. Assuming the double declining balance method of depreciation is used, what is depreciation expense that needs to be recorded at the end of the second year?
Answer #1
Double declining method = (2 * straight line depreciation percent * book value at the beginning)
Straight line depreciation percent = [SLM Depreciaion per year / (Original Cost - Salvage value)]
Straight line depreciation percent = [(64,588 / (565,000-48,300)] = 12.5%
SLM Depreciaion per year = (565,000-48,300)/8 = $64,588
1st Year depreciation = [(2*12.5%*565,000) * 4/12] = $47.083
Book value at the beginning of 2nd Year = (565,000-47,083) = $517,917
2nd year depreciation $[(2*12.5%*517,917)] = $129,479